OpenStudy (anonymous):

Choose the equation of the line passing through the point (1, -5) and perpendicular to y = 1 fifthx - 10. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4 I think that it is B because of how I worked out the problem but I am not sure that I am doing this right. Could someone help me out?

OpenStudy (texaschic101):

y = 1/5x - 10 In y = mx + b form, the slope is in the m position. Therefore, the slope of the given line is 1/5. But we are looking for a perpendicular line, therefore, we need the negative reciprocal of that slope. All that means is " flip " the slope over , and change the sign. So the negative reciprocal of 1/5 is -5/1 or just -5 (see how I flipped the slope and changed the sign). So we need to use -5 as our new slope. y = mx + b slope(m) = -5 (1,-5) x = 1 and y = -5 now we sub and solve for b -5 = -5(1) + b -5 = -5 + b -5 + 5 = b 0 = b your perpendicular equation is : y = -5x + 0 or just y = -5x